OpenStudy (anonymous):

Solve. 2x2 - x = 15 x = 5 over 2 and x = 7 x = 7 over 2 and x = -1 x = 3 and x = - 5 over 2 x = 3 and x = 1

OpenStudy (anonymous):

i'm not sure if i understand what 2x2 means

OpenStudy (anonymous):

2 times 2?

OpenStudy (anonymous):

ohh i get it. 2x^2-x=15 ?

OpenStudy (anonymous):

just let me know if that is correct.

jimthompson5910 (jim_thompson5910):

2x^2-x=15 2x^2-x-15 = 0 that's in the form ax^2 + bx + c = 0 so a = 2 b = -1 c = -15 Plug those into the quadratic formula \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

\[1 \pm \sqrt{4 - 4(2)(-15)} \div 4\]

OpenStudy (anonymous):

you get it..?

OpenStudy (anonymous):

see the formula again...1st term in bracket is b...and b=-1 so b^2 =1 not =4..??

OpenStudy (anonymous):

oh oops heehe

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\Large x=\frac{-(-1) \pm \sqrt{(-1)^2-4(2)(-15)}}{2(2)}\] \[\Large x=\frac{1 \pm \sqrt{1-(-120)}}{4}\] \[\Large x=\frac{1 \pm \sqrt{121}}{4}\] I'll let you finish up

OpenStudy (anonymous):

idk how to simplify that? do you add 1 to 121? do you divide it all by 4??

jimthompson5910 (jim_thompson5910):

first you have to take the square root of 121

OpenStudy (anonymous):

its 11

jimthompson5910 (jim_thompson5910):

So we have \[\Large x=\frac{1 \pm 11}{4}\] which breaks down into these two equations \[\Large x=\frac{1 + 11}{4} \text{ or } x=\frac{1 - 11}{4}\]

jimthompson5910 (jim_thompson5910):

do you see how to go from here?

OpenStudy (anonymous):

yes thank you:)

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

wait no i dont know which answer it is :(

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{1 + 11}{4}=???\]

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{1 - 11}{4} = ???\]

OpenStudy (anonymous):

OHHHHH

jimthompson5910 (jim_thompson5910):

what's that equal to